# note-name = , use-sort-key = false

### v3

@@ Abstract
We investigate the transient dynamics of the quantum Stuart-Landau oscillator, the paradigm of a quantum system exhibiting the quantum limit cycle and synchronization. By considering the energy dynamics, we determine a condition for the validity of the classical-regime approximation—the classical regime eligibility. We also determine the condition for the limit cycle’s classical regime and formulate a guess function that fits the steady-state Wigner function. The equation of motion for the Wigner function is derived and compared to the Kramers-Moyal equation for stochastic processes. We study the classical-like behavior embodied by the coherent state in the quantum phase space, comparing its evolution with or without the classical regime eligibility and attributing its lengthy evolution to the slowness of the neighboring-level coherence decay. Furthermore, we consider the evolution of the Wigner negativity as an indicator of quantumness. The Wigner negativity can increase in some cases due to the nonlinear dissipation. To study the evolution speed, we analyze the Liouvillian gap of the system and record the time to reach the steady state for the Fock, thermal, and coherent states. The steady-state time may behave differently from the Liouvillian gap, depending on the initial state. For the diagonal states, there are speedy parameters for which the steady-state time is significantly lowered.
@@ I Introduction
Introducing some forms of nonlinearities to a simple harmonic oscillator can make it self-sustaining. A self-sustained oscillator is an oscillator whose phase space exhibits one or more stable trajectories to which a phase point is attracted, corresponding to equilibria between the processes influencing it. That is, it can hold its rhythm against transient perturbations. Furthermore, it can synchronize by coupling with another self-sustained oscillator. A self-sustained oscillator is characterized by at least one form of energy gain and energy loss—at least one of which is nonlinear—enabling the equilibrium [1, 2, 3]. The form of the limit cycle depends on the mathematical description of the system [4, 5], and multiple limit cycles can exist for more complex systems [6]. Examples of self-sustaining oscillators include wall clocks, electronic systems [7], the circadian rhythm [8, 9], chemical reactions [10, 11], and predator-and-prey systems [12].
@@
The quantum version of self-sustained oscillators has been gaining interest in recent years. One distinct feature of quantum self-sustained oscillators is the spreading of the limit cycle due to quantum noise [13, 14], a consequence of the uncertainty principle. Typical mathematical descriptions use the Lindblad master equation [15, 16, 17, 18, 14, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], although different formalism is possible [32]. Recent research involving quantum self-sustained oscillators is mostly concerned with quantum synchronization, such as quantum synchronization under external driving [18, 20, 14, 22, 19, 21, 30], squeezing [23, 22], and coupling to other oscillators [24, 25, 26, 27, 28, 30, 31, 29]. Intriguing results were found about quantum frequency or phase entrainment, quantum fluctuations [26, 19], critical response [20], information measures [22], classicality [22, 27, 23], entanglement [25], and amplitude death or oscillation collapse [28, 29, 30]. Several studies also proposed quantifying the degree of quantum synchronization [33, 34, 35, 36]. Furthermore, research has been done to speed up synchronization via drive engineering [37]. Apart from quantum synchronization, quantum limit cycles also play an important role in heat engines [15, 16] and single-electron transistors [17].
@@
Studies have also been done on the dynamics of sole quantum self-sustained oscillators. However, they are mainly concerned with the steady-state dynamics—the quantum limit cycles—apart from the foundational work of quantizing the equations of motion [13, 14]. The full dynamics toward the quantum limit cycle are much less explored. Unlike the classical case, the motion of the phase point attracted to the limit cycle is not the only information to obtain. As such, we may anticipate seeing interesting features of the self-sustaining system. These features may additionally be useful for solving related problems, such as those in quantum synchronization.
In this work, we study the transient dynamics of the quantum Stuart-Landau (SL) oscillator, the prime example of a quantum self-sustained oscillator exhibiting the quantum limit cycle and synchronization. We describe the absence of the system’s phase point in the classical limit, which is attributed to quantum decoherence. By considering how the system’s energy evolves, the conditions for the validity of the classical-regime approximation—which we call the classical-regime eligibility—and the limit cycle’s classical regime can be formulated. The equation of motion for the Wigner function is derived and its form is compared to the Kramers-Moyal (KM) equation for stochastic processes. We then study the system’s evolution starting with the classical-like coherent state, comparing the evolution as the system gains or loses the classical regime eligibility. Interestingly, the decay of the coherence between eigenstates of neighboring energy levels is slower than the radial attraction.
@@
We also consider the negative region of the Wigner function, which embodies a state’s quantumness, where we find that the Wigner negativity can increase early in the evolution due to the two-quantum dissipation for some initial states. Moreover, we numerically calculate the Liouvillian gap and study its dependence on the system’s parameters to study the speed of system’s evolution. Lastly, we evolve the system from different initial states and record the time to reach steady-state for several varying parameters. We find that, generally, the evolution does not behave like the Liouvillian gap. For the coherent state, the steady-state time behaves similarly to the Liouvillian gap and does not depend on its initial radial position in phase space. Meanwhile, the steady-state time significantly drops for specific sets of parameters. These speedy parameters are observed for the Fock and coherent states, suggesting that they are attributed to the system’s energy balance.
@@
The rest of this paper is organized as follows. In Section II, we introduce the mathematical model of the system and discuss some general properties: the connection to the classical SL oscillator, the quantum limit cycle, the disappearance of the phase point due to decoherence, the energy dynamics, and the classical regime. The equation of motion for the system’s Wigner function, the features of the evolution of the classical-like coherent state, and the evolution of the Wigner negativity are discussed in Section III. We consider the Liouvillian gap and characterize the steady-state time for selected initial states in Section IV. Finally, Section V presents the conclusions and outlooks of our study.
@@ II The System
We consider a quantum self-sustained oscillator described by the following Lindblad master equation:
$\begin{split}\rho^{\prime}&=\mathcal{L}\rho\end{split}$ (1)
Here the Liouville superoperator in Lindblad form (or simply Lindbladian) $\mathcal{L}$ is given by
$\begin{split}\mathcal{L}\rho&=-i\left[{\hat{H}_{0}},{\rho}\right]\\ &\quad+\kappa_{1}\mathcal{D}\left({\hat{a}^{\dagger}}\right)\left[\rho\right]+\gamma_{1}\mathcal{D}\left({\hat{a}}\right)\left[\rho\right]+\gamma_{2}\mathcal{D}\left({\hat{a}^{2}}\right)\left[\rho\right]\end{split}$ (2)
where $\rho^{\prime}=\mathrm{d}\rho/\mathrm{d}t$, $\hat{H}_{0}=\hat{a}^{\dagger}\hat{a}+1/2$, and $\mathcal{D}\left({\hat{O}}\right)\left[\rho\right]=\hat{O}\rho\hat{O}^{\dagger}-(1/2)\left\{{\hat{O}^{\dagger}\hat{O}},{\rho}\right\}$. In Eq. (1), inverse time is measured in units of the oscillator’s natural frequency $\omega_{0}$ (making the simple harmonic period $2\pi$), mass in units of the oscillator’s mass $m$, and length in units of $\sqrt{\hbar/m\omega_{0}}$. This unit system follows Ref. [13] and makes Eq. (1) dimensionless. Along with $\hbar=1$, the ladder operators $\hat{a}=\left(m\omega_{0}\hat{x}+i\hat{p}\right)/\sqrt{2\hbar m\omega_{0}}$ becomes simply $\hat{a}=\left(\hat{x}+i\hat{p}\right)/\sqrt{2}$ in this unit system. There are three sources of non-conservative mechanisms responsible for the self-sustaining property: $\kappa_{1}$ parameterizes the rate of one-quantum pump, $\gamma_{1}$ parameterizes the rate of one-quantum dissipation, and $\gamma_{2}$ parameterizes the rate of two-quantum dissipation. This master equation has been used in the literature, such as by Refs. [13, 20, 38], and is a generalization of another model without the $\gamma_{1}$ process widely used in the topics of limit cycles and synchronization [18, 14, 19, 23, 24, 25, 26, 28, 31]. Experimental realization of the system in optomechanical setups [18, 24, 39], trapped-ion systems [26], and spin systems [40] have been proposed.
@@ II.1 The system and the classical Stuart-Landau oscillator
The expected phase point dynamics of the system is given by the time dependence of $\left\langle{\hat{a}}\right\rangle=\mathrm{tr}\left({\rho\hat{a}}\right)$. Operating Eq. (1) by $\hat{a}$ from either side and using the commutation relation $\left[{\hat{a}},{\hat{a}^{\dagger}}\right]=1$, we obtain
$\left\langle{\hat{a}}\right\rangle^{\prime}=-i\left\langle{\hat{a}}\right\rangle+\frac{\kappa_{1}-\gamma_{1}}{2}\left\langle{\hat{a}}\right\rangle-\gamma_{2}\left\langle{\hat{a}^{\dagger}\hat{a}^{2}}\right\rangle$ (3)
@@
$\alpha^{\prime}_{\mathrm{c}}=-i\alpha_{\mathrm{c}}+\frac{\kappa_{1}-\gamma_{1}}{2}\alpha_{\mathrm{c}}-\gamma_{2}|\alpha_{\mathrm{c}}|^{2}\alpha_{\mathrm{c}}$ (4)
@@
$r^{\prime}_{\mathrm{c}}$ $=\frac{(\kappa_{1}-\gamma_{1})r_{\mathrm{c}}-\gamma_{2}r_{\mathrm{c}}^{3}}{2}$ (5a)
$\theta^{\prime}_{\mathrm{c}}$ $=-1$ (5b)
The phase point rotates clockwise uniformly at the simple harmonic frequency while being radially attracted toward the circular limit cycle of radius $r_{\mathrm{c,lc}}=\sqrt{(\kappa_{1}-\gamma_{1})/\gamma_{2}}$bibnoteWhen $\kappa_{1}<\gamma_{1}$, the limit cycle radius $r_{\mathrm{c,lc}}$ becomes imaginary and is thus unphysical. Since the energy pump loses to the linear damping, the two terms in Eq. (5a) synergize in decreasing $r_{\mathrm{c}}$ and the oscillation dies out., obtained by setting the LHS of Eq. (5a) to zero and solving for $r_{\mathrm{c}}$. If $r_{\mathrm{c}}<r_{\mathrm{c,lc}}$, the term $(\kappa_{1}-\gamma_{1})r_{\mathrm{c}}$ is larger so $r_{\mathrm{c}}$ increases and vice versa. This corresponds to the SL oscillator’s simple processes to self-sustain its oscillation: (i) a net energy pump when the energy is below the limit-cycle energy and (ii) a net energy dissipation otherwise.
@@ II.2 The quantum limit cycle
The limit cycle of a quantum oscillator is exhibited in the phase space representation of the steady-state density matrix [13, 14], analogous to how the limit cycle of a classical oscillator is exhibited in its phase space. In the quantum phase-space representation, we have a quasiprobability distribution that represents the density matrix and—like the density matrix—does not tell us much about the system’s observables without being treated under the appropriate mathematical operations [42, 43, 44]. This sets the quantum limit cycle apart from the classical counterpart—the quantum limit cycle is generally not just the classical limit cycle spread out into a probability distribution due to the quantum noise. Nevertheless, by looking at the Wigner function, we may discern some properties of the quantum state that may not be directly evident from the density matrix—for example, negative regions guaranteeing nonclassicality, entanglement, the expected phase point of a coherent state, and squeezing of the quantum uncertainty.
Ref. [13] has previously described the transient dynamics in terms of the density matrix $\rho$ and thoroughly studied the steady-state dynamics. Let the “$m$th off-diagonal” be the density matrix entries $\rho_{n,n+m}$. A given density matrix entry is only coupled to other entries along the same off-diagonal. None of the off-diagonal entries are coupled to the diagonal $(m=0)$ elements and decay to zero, showing that any coherence initially present in the system disappears along the evolution. Meanwhile, the occupation probabilities $P_{n}=\left\langle{n}\middle|{\rho}\middle|{n}\right\rangle$ on the density matrix diagonal redistributes toward the steady-state distribution, given by
$P_{n}^{\mathrm{(ss)}}=\frac{\tilde{\kappa}_{1}^{n}}{\left(\tilde{\kappa}_{1}+\tilde{\gamma}_{1}\right)_{n}}\frac{\prescript{}{1}{F_{1}}(1+n;\tilde{\kappa}_{1}+\tilde{\gamma}_{1}+n;\tilde{\kappa}_{1})}{\prescript{}{1}{F_{1}}(1;\tilde{\kappa}_{1}+\tilde{\gamma}_{1};2\tilde{\kappa}_{1})}$ (6)
where $(x)_{n}=x(x+1)(x+2)\dots(x+n-1);\ (x)_{0}=1$ is the Pochhammer’s symbol, $\prescript{}{1}{F_{1}}(b;c;z)=\sum_{n=0}^{\infty}[(b)_{n}z^{n}]/[(c)_{n}n!]$ is the confluent hypergeometric function [45], $\tilde{\kappa}_{1}=\kappa_{1}/\gamma_{2}$, and $\tilde{\gamma}_{1}=\gamma_{1}/\gamma_{2}$. The steady state $\rho^{\mathrm{(ss)}}$—obtained by solving $\mathcal{L}\rho^{\mathrm{(ss)}}=0$—is a statistical mixture of pure states.
@@
$W(x,p)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\left\langle{x-\frac{y}{2}}\middle|{\rho}\middle|{x+\frac{y}{2}}\right\rangle e^{ipy}\ \mathrm{d}y$ (7)
Since $\rho^{\mathrm{(ss)}}$ is diagonal, the corresponding Wigner function $W^{\mathrm{(ss)}}$ is rotationally invariant in the phase space, characterizing the quasiharmonic nature of the limit cycle. Ref. [13] has previously described some interesting properties of $W^{\mathrm{(ss)}}$. In the regime where $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}=\mathrm{tr}\left({\rho^{\mathrm{(ss)}}\hat{a}^{\dagger}\hat{a}}\right)\gg 1$, the circular peak of $W^{\mathrm{(ss)}}$ coincides with the classical limit cycle amplitude of the corresponding SL oscillator. As $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ becomes smaller, the peak shrinks in radius, deviating from the classical SL limit cycle. The peak’s radius saturates near the zero-point radius $r_{\mathrm{zp}}=1/\sqrt{2}$ as $\gamma_{2}\rightarrow\infty$, after which point the radius further decreases toward zero as the ratio $\gamma_{1}/\kappa_{1}$ gets closer to unity. The Wigner function exhibits Hopf bifurcation (i.e. the boundary between the system’s limit cycle and damping regime [2]) at $\kappa_{1}=\gamma_{1}$, similar to the classical SL oscillator.

### v5

# Transient dynamics of the quantum Stuart-Landau oscillator

@@ Abstract
We investigate the transient dynamics of the quantum Stuart-Landau oscillator, a paradigmatic quantum system exhibiting a quantum limit cycle and synchronization. From the energy dynamics, we determine a condition for the classical regime of transient dynamics and the limit cycle. Additionally, we formulate a guess function that fits the classical-regime steady-state Wigner function. The equation of motion for the Wigner function is derived and compared to the Kramers-Moyal equation for stochastic processes. We then characterize the classical-like behavior as the system evolves from a coherent state, noting the slow decay of neighboring-level coherence. We also study the evolution of the Wigner negativity as an indicator of nonclassicality, showing its temporary increase for some specific cases. To quantify the evolution speed, we examined the system’s Lindbladian spectra, particularly the Liouvillian gap. Finally, we record the time it takes to reach the steady state for some Fock, thermal, and coherent states. The parameter dependence of the steady-state time may differ from the Liouvillian gap, and the limit-cycle attraction is significantly slower for coherent states compared to Fock or thermal states. For the diagonal states, there are fast convergence regimes for which the steady-state time is locally minimized. This study provides a deeper insight into the transient behavior of self-sustained quantum systems.
@@ I Introduction
Introducing nonlinearities to a simple harmonic oscillator can make it self-sustaining. A self-sustained oscillator is a dynamical system whose phase space exhibits one or more stable trajectories to which a phase point is attracted, corresponding to equilibria between the processes influencing it. It can hold its rhythm against transient perturbations and synchronize via coupling with one another. A self-sustained oscillator is characterized by at least one form of energy gain and energy loss (at least one of which is nonlinear), enabling the equilibrium [1, 2, 3]. The form of the limit cycle depends on the mathematical description of the system [4, 5], and multiple limit cycles can exist for more complex systems [6]. Examples of self-sustaining oscillators include wall clocks, electronic systems [7], the circadian rhythm [8, 9], chemical reactions [10, 11], and predator-and-prey systems [12].
@@
The quantum version of self-sustained oscillators has been gaining interest in recent years. One distinct feature of quantum self-sustained oscillators is the spreading of the limit cycle due to quantum noise [13, 14], a consequence of the uncertainty principle. Typical mathematical descriptions use the Lindblad master equation [15, 16, 17, 18, 14, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], although different formalisms may be possible [32]. Recent research on quantum self-sustained oscillators is mostly concerned with quantum synchronization, such as quantum synchronization under external driving [18, 20, 14, 22, 19, 21, 30], squeezing [23, 22], and coupling to other oscillators [24, 25, 26, 27, 28, 30, 31, 29]. Intriguing results regarding quantum frequency or phase entrainment, quantum fluctuations [26, 19], critical response [20], information measures [22], classicality [22, 27, 23], entanglement [25], and amplitude death or oscillation collapse [28, 29, 30] have been reported. Several studies have also proposed quantifying the degree of quantum synchronization [33, 34, 35, 36], and to speed up synchronization via drive engineering [37]. In addition to quantum synchronization, quantum self-sustained oscillators also play an important role in heat engines [15, 16] and single-electron transistors [17].
@@
Studies have also been done on the dynamics of sole quantum self-sustained oscillators. However, they are mainly concerned with steady-state dynamics, that is, quantum limit cycles, apart from the foundational work of quantizing the equations of motion [13, 14]. The transient dynamics toward the quantum limit cycle is much less explored. Recently, Ref. [38] studies how a quantum limit cycle may generally emerge, but we believe there is more to uncover. Unlike a classical phase space, richer phenomena like interference, decoherence, and nonclassicality are observable in the quantum phase space. We may anticipate interesting features that are potentially useful for solving related problems, such as those in quantum synchronization.
In this work, we study the transient dynamics of the quantum Stuart-Landau (SL) oscillator, the paradigm of a quantum self-sustained oscillator exhibiting quantum limit cycle and synchronization. By considering the system’s energy evolution, the conditions for the validity of the classical-regime approximation, i.e., the classical-regime eligibility, and the limit cycle’s classical regime can be formulated. We also propose a guess function for the classical-regime steady-state Wigner function, which exhibits minimal quantum noise signifying the approach toward the classical limit.
Moving onto the quantum phase-space representation of the transient dynamics, we derive the equation of motion for the Wigner function. We relate the resulting terms to the features of the Wigner function evolution by comparing them with the Kramers-Moyal (KM) equation for stochastic processes, revealing some deviations due to the quantum nature. We then study the system’s evolution starting from a classical-like coherent state and observe how the dynamics vary as the system gains or loses the classical regime eligibility. Additionally, we find that the decay of coherence between the eigenstates of neighboring energy levels is slower than the radial attraction corresponding to the eigenstate occupation probability redistribution. We also consider the negative region of the Wigner function, which signifies nonclassicality. For some initial states, we observe that the Wigner negativity may temporarily increase during the evolution; this effect is attributed to the two-quantum dissipation.
@@
To characterize the evolution speed, we numerically evaluate the system’s Lindbladian spectrum and study its dependence on the system’s parameters. We specifically consider the Liouvillian gap, which governs the slowest dynamics toward the limit cycle, noting the false turning point that arises due to bad Hilbert space truncation in the procedure. Lastly, we evolve the system from different initial states (coherent, Fock, thermal) and record the time it takes to reach the steady state. We find that, generally, the evolution does not behave like the Liouvillian gap. For the coherent state, the steady-state time behaves similarly to the Liouvillian gap and does not depend on its initial radial position in phase space. The system takes significantly more time to reach the limit cycle starting as a coherent state than the diagonal states (Fock, thermal), which we attribute to the slow decay of neighboring-level coherence. We observe that the steady-state time is locally minimized over some parameter ranges; these fast convergence regimes are observed for the diagonal states, suggesting that they are attributed to the system’s energy balance.
@@
The remainder of this paper is organized as follows. In Section II, we introduce the mathematical model of the system and discuss some general properties: the connection to the classical SL oscillator, the quantum limit cycle, the disappearance of the phase point due to decoherence, the energy dynamics, and the classical regime. The equation of motion for the system’s Wigner function, the features of the evolution of the classical-like coherent state, and the evolution of the Wigner negativity are discussed in Section III. We consider the Liouvillian gap and characterize the steady-state time for selected initial states in Section IV. Finally, Section V presents the conclusions and perspectives of our study.
@@ II The System
We consider a quantum self-sustained oscillator described by the following Lindblad master equation,
$\begin{split}\rho^{\prime}&=\mathcal{L}\rho.\end{split}$ (1)
Here, the Liouville superoperator in Lindblad form (or simply Lindbladian) $\mathcal{L}$ is given by
$\begin{split}\mathcal{L}\rho&=-i\left[{\hat{H}_{0}},{\rho}\right]\\ &\quad+\kappa_{1}\mathcal{D}\left({\hat{a}^{\dagger}}\right)\left[\rho\right]+\gamma_{1}\mathcal{D}\left({\hat{a}}\right)\left[\rho\right]+\gamma_{2}\mathcal{D}\left({\hat{a}^{2}}\right)\left[\rho\right],\end{split}$ (2)
where $\rho^{\prime}=\mathrm{d}\rho/\mathrm{d}t$, $\hat{H}_{0}=\hat{a}^{\dagger}\hat{a}+1/2$, and $\mathcal{D}\left({\hat{O}}\right)\left[\rho\right]=\hat{O}\rho\hat{O}^{\dagger}-(1/2)\left\{{\hat{O}^{\dagger}\hat{O}},{\rho}\right\}$. In Eq. (1), inverse time is measured in units of the oscillator’s natural frequency $\omega_{0}$ (making the simple harmonic period $2\pi$), mass in units of the oscillator’s mass $m$, and length in units of $\sqrt{\hbar/m\omega_{0}}$. This unit system follows Ref. [13] and makes Eq. (1) dimensionless. Along with $\hbar=1$, the ladder operators $\hat{a}=\left(m\omega_{0}\hat{x}+i\hat{p}\right)/\sqrt{2\hbar m\omega_{0}}$ becomes simply $\hat{a}=\left(\hat{x}+i\hat{p}\right)/\sqrt{2}$ in this unit system. There are three sources of nonconservative mechanisms responsible for the self-sustaining property: $\kappa_{1}$ parameterizes the rate of one-quantum pump, $\gamma_{1}$ parameterizes the rate of one-quantum dissipation, and $\gamma_{2}$ parameterizes the rate of two-quantum dissipation. This master equation has been used in the literature, e.g., in Refs. [13, 20, 39], and is a generalization of another model without the $\gamma_{1}$ process widely used in the topics of limit cycles and synchronization [18, 14, 19, 23, 24, 25, 26, 28, 31]. In particular, the physical realizations of the nonconservative processes are discussed in Section III A in Ref. [13]. Experimental realization of the system in optomechanical setups [18, 24, 40], trapped-ion systems [26], and spin systems [41] have been proposed.
@@ II.1 The system and the classical Stuart-Landau oscillator
[…]
$\left\langle{\hat{a}}\right\rangle^{\prime}=-i\left\langle{\hat{a}}\right\rangle+\frac{\kappa_{1}-\gamma_{1}}{2}\left\langle{\hat{a}}\right\rangle-\gamma_{2}\left\langle{\hat{a}^{\dagger}\hat{a}^{2}}\right\rangle.$ (3)
@@
$\alpha^{\prime}_{\mathrm{c}}=-i\alpha_{\mathrm{c}}+\frac{\kappa_{1}-\gamma_{1}}{2}\alpha_{\mathrm{c}}-\gamma_{2}|\alpha_{\mathrm{c}}|^{2}\alpha_{\mathrm{c}},$ (4)
@@
$r^{\prime}_{\mathrm{c}}$ $=\frac{(\kappa_{1}-\gamma_{1})r_{\mathrm{c}}-\gamma_{2}r_{\mathrm{c}}^{3}}{2},$ (5a)
$\theta^{\prime}_{\mathrm{c}}$ $=-1.$ (5b)
The phase point rotates clockwise uniformly at the simple harmonic frequency while being radially attracted toward the circular limit cycle of radius $r_{\mathrm{c,lc}}=\sqrt{(\kappa_{1}-\gamma_{1})/\gamma_{2}}$  [^1] When $\kappa_{1}<\gamma_{1}$, the limit cycle radius $r_{\mathrm{c,lc}}$ becomes imaginary and is thus unphysical. Since the energy pump loses to the linear damping, the two terms in Eq. (5a) synergize in decreasing $r_{\mathrm{c}}$ and the oscillation dies out., obtained by setting the LHS of Eq. (5a) to zero and solving for $r_{\mathrm{c}}$. If $r_{\mathrm{c}}<r_{\mathrm{c,lc}}$, the term $(\kappa_{1}-\gamma_{1})r_{\mathrm{c}}$ is larger so $r_{\mathrm{c}}$ increases and vice versa. This corresponds to the SL oscillator’s simple processes to self-sustain its oscillation: (i) a net energy pump when the energy is below the limit-cycle energy, and (ii) a net energy dissipation otherwise.
@@ II.2 The quantum limit cycle
The limit cycle of a quantum oscillator is exhibited in the phase space representation of the steady-state density matrix [13, 14], analogous to how the limit cycle of a classical oscillator is exhibited in its phase space. In the quantum phase-space representation, we have a quasiprobability distribution that represents the density matrix and, like the density matrix, does not tell us much about the system’s observables without being treated under the appropriate mathematical operations [43, 44, 45]. This sets the quantum limit cycle apart from the classical counterpart: the quantum limit cycle is generally not just the classical limit cycle spread out into a probability distribution due to the quantum noise. Nevertheless, by observing the Wigner function, we can discern properties of the quantum state that may not be directly evident from the density matrix, e.g., negative regions indicating nonclassicality, entanglement, the expected phase point of a coherent state, and squeezing of the quantum uncertainty.
Ref. [13] has previously described the transient dynamics in terms of the density matrix $\rho$ and thoroughly studied the steady-state dynamics. Let the $m$th off-diagonal be the density matrix entries $\rho_{n,n+m}$. A given density matrix entry is only coupled to other entries along the same off-diagonal. None of the off-diagonal entries are coupled to the diagonal $(m=0)$ elements and decay to zero, showing that any coherence initially present in the system disappears along the evolution. Meanwhile, the occupation probabilities $P_{n}=\left\langle{n}\middle|{\rho}\middle|{n}\right\rangle$ on the density matrix diagonal redistribute toward the steady-state distribution, given by
$P_{n}^{\mathrm{(ss)}}=\frac{\tilde{\kappa}_{1}^{n}}{\left(\tilde{\kappa}_{1}+\tilde{\gamma}_{1}\right)^{\overline{n}}}\frac{\prescript{}{1}{F_{1}}(1+n;\tilde{\kappa}_{1}+\tilde{\gamma}_{1}+n;\tilde{\kappa}_{1})}{\prescript{}{1}{F_{1}}(1;\tilde{\kappa}_{1}+\tilde{\gamma}_{1};2\tilde{\kappa}_{1})},$ (6)
where $(x)^{\overline{n}}=x(x+1)(x+2)\dots(x+n-1);\ (x)^{\overline{0}}=1$ is Pochhammer’s symbol, $\prescript{}{1}{F_{1}}(b;c;z)=\sum_{n=0}^{\infty}[(b)^{\overline{n}}z^{n}]/[(c)^{\overline{n}}n!]$ is the confluent hypergeometric function [46], $\tilde{\kappa}_{1}=\kappa_{1}/\gamma_{2}$, and $\tilde{\gamma}_{1}=\gamma_{1}/\gamma_{2}$. The steady state $\rho^{\mathrm{(ss)}}$, obtained by solving $\mathcal{L}\rho^{\mathrm{(ss)}}=0$, is a statistical mixture of pure states.
@@
$W(x,p)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\left\langle{x-\frac{y}{2}}\middle|{\rho}\middle|{x+\frac{y}{2}}\right\rangle e^{ipy}\ \mathrm{d}y.$ (7)
Since $\rho^{\mathrm{(ss)}}$ is diagonal, the corresponding Wigner function $W^{\mathrm{(ss)}}$ is rotationally invariant in the phase space, characterizing the quasiharmonic nature of the limit cycle. Ref. [13] has previously described some interesting properties of $W^{\mathrm{(ss)}}$. In the regime where $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}=\mathrm{tr}\left({\rho^{\mathrm{(ss)}}\hat{a}^{\dagger}\hat{a}}\right)\gg 1$, the circular peak of $W^{\mathrm{(ss)}}$ coincides with the classical limit cycle amplitude of the corresponding SL oscillator. As $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ becomes smaller, the peak shrinks in radius, deviating from the classical SL limit cycle. The peak’s radius saturates near the zero-point radius $r_{\mathrm{zp}}=1/\sqrt{2}$ as $\gamma_{2}\rightarrow\infty$, after which the radius further decreases toward zero as the ratio $\gamma_{1}/\kappa_{1}$ gets closer to unity. The Wigner function exhibits Hopf bifurcation (i.e., the boundary between the system’s limit cycle and damping regime [2]) at $\kappa_{1}=\gamma_{1}$, similar to the classical SL oscillator.